OpenStudy (readergirl12):

Solve the equation for y. 8x – 9y = 11 A. y= -8/9x - 9/11 B. y= 9/8x + 11/8 C. y= -8x - 11 D. y= 8/9x - 11/9

pooja195 (pooja195):

We need to solve for y so start by subtracting 8x from both sides \[\huge~\rm~8x -8x– 9y = 11-8x\] What equation are u left with?

OpenStudy (igigighjkl):

X= Y= if you sub the two numbers to were they both equal 11, you get the awnser.

OpenStudy (readergirl12):

D? v-v

OpenStudy (igigighjkl):

what does that mean?

pooja195 (pooja195):

Incorrect

OpenStudy (igigighjkl):

I'm incorrect? aw.

OpenStudy (calculusxy):

This is in standard form: \(\large ax + by = c\) And the slope-intercept form is \(\large y = mx + b\). Thus, you need to have \(\large y\) on one side and \(\large ax\) and \(\large c\) on the other (note that \(\large ax\) sort of means \(mx\) in the slope-int form and \(\large c\) means \(\large b\) in slope-int form).

pooja195 (pooja195):

8x-8x cancels out now you have -9y=-8x+11 divide everything by -9 \[\huge~\rm~\bf-9y/-9=-8x/-9+11 /-9\]

OpenStudy (igigighjkl):

O.k i'm lost.

OpenStudy (readergirl12):

20 and 0?

pooja195 (pooja195):

-9/-9=1 so just y so now you have y=-8x/-9+11/-9

OpenStudy (igigighjkl):

sorry can't help cause i don't know this.

OpenStudy (igigighjkl):

sorry can't help cause i don't know this.

OpenStudy (calculusxy):

Similar to what @pooja195 is doing... \(\large 8x - 9y = 11\) \(\large 8x - 8x - 9y = 11 - 8x\) \(-9y = 11 - 8x\) divide both sides by \(-9\) to isolate \(y\). \(y = -\frac{11}{9} + \frac{8}{9}x\) we can now rearrange this for a better interpretation. \(\large y = \frac{8}{9}x - \frac{11}{9}\)

OpenStudy (readergirl12):

Thank you peeps.. xDD

OpenStudy (calculusxy):

You're welcome :)